OpenStudy (anonymous):

A 1230kg car pushes a 2160kg truck that has a dead battery. When the driver steps on the accelerator, the drive wheels of the car push against the ground with a force of 4590N . Take the friction between the truck's tires and the ground to be 470N . What is the magnitude F_ConT of the force that the car exerts on the truck? Use the coordinate system shown in Part B, where upward is the positive y direction and to the right, the direction F_ConT points, is the positive x direction.

OpenStudy (anonymous):

i'm on it

OpenStudy (anonymous):

the value of F for c on t is 2796N

OpenStudy (anonymous):

That answer is incorrect. The value I got was close to that one.

OpenStudy (anonymous):

when yu find the accelelration, it solves everything

OpenStudy (anonymous):

\[a = \frac{ (F _{ConT}-470) }{ 2160 }, a = \frac{ 4590-F _{TonC} }{ 1230 }\]

OpenStudy (anonymous):

\[a=\frac{ F-2f _{f} }{ m _{c} +m _{t}}\] \[F-f _{f}-f _{t on c} = m _{c}a\] \[f _{c on t} - f _{f} = m _{t}a\]

OpenStudy (anonymous):

so if if yu add the two equations, the f t on c cancels with the f c on t, cause they are the same just different direction

OpenStudy (anonymous):

What is capital F in your equation.

OpenStudy (anonymous):

I think your setup may be correct but the two friction forces are different in this environment

OpenStudy (anonymous):

the F is the N no then we will have to use \[muN\] as the friction force

OpenStudy (anonymous):

Got the answer, thank you.

OpenStudy (anonymous):

which did u use? which formula?

OpenStudy (anonymous):

\[a = \frac{f _{C}-f _{T} }{ M _{C}+M _{T}} \rightarrow M _{T}(a)+f _{T}=F _{ConT}\]